OpenStudy (bbb911):

Multiply to lowest term 16/21 X 3/5 X 5/8

OpenStudy (anonymous):

\[\frac{ 16 }{ 21 } \times \frac{ 3 }{ 5 } \times \frac{ 5 }{ 8 } = \frac{ 240 }{ 840 } =\frac{ 2 }{ 7 }\]

OpenStudy (bbb911):

thank you so much :) :)